2 The table below represents how Marco feels about chocolate candy bars. a. Fill in the missing values for total utility and marginal utility Instructions: Enter your answers as a whole number. Chocolate Candy Bars and Marcos Utility 0.17 points Chocolate Candy Bars Total Utility Marginal Utility (utils) (utils) 29 Print 2 21 4. 86 8 6 Suppose Marco currently has two candy bars. You tell Marco you will give him either a soda, which gives him 29 utils 0hpp ness,or two additional candy bars. b. Which is he likely to prefer? (Click to select
Answer #1

Question 2

A)

MUn = TUn + TUn-1

Here,

  • MUn represents Marginal Utility From the nth Unit

  • TUn represents Total Utility From n Units

  • TUn-1 represents Total Utility From n-1 units

  • n represents number of units of consumption

Using the above formala we can calculate the Total Utility and the Marginal Utility  
For Example: MU1 = TU1 + TU0
MU1 = 29 - 0 = 29 Utilis

Likewise Solving all the unfilled areas in the table we the below table (highlighted with Yellow Colour)Chocolete Candylul Cutii) 29 291 3 S: T2 2 g95

B) Macros has two candy bar with him, on giving two additional candy bar, he will be having in total 4 candy bars
Total Utility Derive from Four Candy Bar is 78 Utils(From the Table), and
Total Utility Derive from Soda is 29 Utils

So Macros will prefer two additional candy bar because additional two candy bar is providing more utils(78 Utils) in comparision to Soda(29 Utils)

Hope This Helps...!!!


answered by: ANURANJAN SARSAM